Operations Report – May 13th

Crew 212 Operations Report 13-MAY-2019

SOL: 8

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Rover Opportunity, Upper deck water pump, Kitchen sink.

Notes on non-nominal systems: rover Opportunity failed during EVA (on the way to the second EVA objective). For more details, look at the notes about the rovers and the EVA report for EVA #9.

The upper deck pump is now activating for rather long times (tens of seconds) with all the sinks shut. The kitchen sink is leaking.

Generator (hours run last night): 11:15

On (last night): 7:45 PM

Off (this morning): 7:00 AM

Solar— SOC 61% at 7:45 PM

Diesel Reading – 88 %

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 380 gallons

Axillary to Static tank transfer – no

Gallons transferred: ~200

Water in GreenHab – 137 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143583.1

Toilet tank emptied: No

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.8h

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.5h

Beginning charge: 100

Ending charge: 98

Currently charging: Yes

Curiosity rover used: Yes

Hours: 95.6h

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: Opportunity rover failed during EVA 9 because of overheating of battery (more details in EVA report); it was not possible to restart the engine until . Problem might be cause by lack of distilled water in radiator. Spirit rover battery has had anomalous consumption rates in the past days. A major rover check will be run tomorrow during an Engineering EVA.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? The HabCar was used to reach the location were Opportunity was left after it failed during EVA 9.

CrewCar used and why, where? Off Campus

General notes and comments: We captured another mouse last night, and it was set free today during EVA 9 at Cowboy’s Corner. EVA 9 presented several challenges (see EVA report).

Summary of internet: Nothing to report.

Summary of suits and radios: some suits batteries are at end of life, they lose charge very quickly (see spacesuit engineer report).

Summary of Hab operations: The ceiling in the commander’s room is compromised by water and might have a localized collapse if weight is applied to it, a bucket had to be placed close to the bed. A mouse was found in the lower deck. Leaking of the sink in the kitchen has been found, and this could be the cause of the leaking reported in a previous report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: what is the level of usage of the potable water filter? It is quite brown at the moment (see photo attached). Should it be replaced together with the pump tomorrow night?

Copyright © The Mars Society. All rights reserved. | Main Site